# Sliding DFT-based Signal Recovery for Modulo ADC with 1-bit Folding Information

### v3

@@ Abstract
The modulo analog-to-digital converter (ADC) is a promising solution to resolve the limited dynamic range (DR) issue of conventional ADCs and achieve an enhanced digital resolution given a fixed quantization bit budget. However, a modulo ADC requires an unfolding scheme to correct the nonlinear distortion introduced by the modulo operation. This paper presents a sliding discrete Fourier Transform (DFT)-based method for fast signal reconstruction given the modulo ADC output sequence and a 1-bit folding information sequence. In contrast to existing DFT-based signal recovery techniques for modulo ADCs, our proposed sliding DFT method reduces the required observation time and minimizes the spectral leakage effects via proper choice of window function parameters. A mean squared error (MSE) performance guarantee is established for the proposed signal recovery algorithm. More precisely, we derive sufficient conditions for the oversampling factor ($\mathrm{OF}$) and the number of quantization bits ($b$) to obtain a specific MSE performance. Our numerical results demonstrate that modulo ADCs equipped with our proposed recovery method can outperform conventional ADCs without modulo for $\mathrm{OF}\geq 4$ and $b\geq 4$. Moreover, the proposed sliding DFT-based recovery method outperforms higher-order difference-based recovery in terms of MSE, particularly in the low-sampling-rate and low-resolution regime. The impact of spectral leakage on the MSE performance of the proposed sliding DFT recovery method is also quantified.
Index Terms:
Modulo ADC, Sampling, Quantization, Sliding DFT, Reconstruction.
@@ I Introduction
An analog-to-digital converter (ADC) is a critical element of data acquisition systems since it is responsible for the transformation of continuous-time analog observations from the physical world into some format that can be further processed or stored in digital signal processing (DSP) pipelines. The conversion process in conventional ADCs involves two operations: (1) a sampler that acquires the amplitude of the continuous-time input at uniform time intervals, and (2) a quantizer that maps the amplitude samples into a finite set of discrete values [1]. The sampling rate and quantizer resolution of an ADC are chosen according to some fidelity criterion and power consumption budget. Theoretically, the sampling rate requirement for alias-free reconstruction and the quantization resolution needed to achieve some distortion level are governed by the Shannon-Nyquist sampling theorem [2] and the rate-distortion theory [3], respectively. However, increasing these parameters also incurs a substantial increase in the ADC power consumption. More precisely, the ADC power consumption scales exponentially with the number of quantization bits and scales linearly with the sampling rate [4].
@@
Another important parameter that affects the performance of an ADC is the dynamic range (DR). Due to the finite number of quantization bits in an ADC, a trade-off exists between its dynamic range and its resolution. An input signal may drive the ADC into saturation if the DR of the input signal far exceeds the DR of the ADC. However, increasing the ADC range to accommodate high DR signals would increase the quantization step size. Weak input signals could be buried in the quantization noise. The presence of both weak and strong components in the input signal increases the required ADC resolution to achieve a specific fidelity criterion [5].
@@
A promising solution to address the DR bottleneck of ADCs is the modulo sampling framework [6, 7]. The idea in modulo sampling is to apply a nonlinear folding operation called modulo to the input signal before the ADC. This modulo pre-processing stage ensures that the signal range is confined within the DR of the ADC. Moreover, for a given amplitude quantization bit budget, modulo ADCs can achieve better digital resolution compared to conventional ADCs as demonstrated in prior works [8, 9].
@@
Since modulo operation intentionally introduces nonlinear distortion to the signal, several studies on modulo sampling have focused on the development of robust recovery techniques to correct the distortion induced by the modulo operation. Reconstruction techniques for modulo sampling include higher-order differences-based approaches [7], discrete Fourier transform (DFT)-based techniques [10, 11, 12, 13, 14, 15], prediction-based filtering [16, 17], iterative methods [18], sparse signal modulo recovery [19, 20, 21, 22], multi-channel modulo structures [23, 24, 25, 26], and thresholding-based schemes[27, 28, 29, 30, 31]. There are also research efforts geared toward the practical hardware implementation of modulo ADCs [10, 32, 33, 34, 35, 36] and their potential applications in radars, communication systems, imaging, and electroencephalogram (EEG) recovery [37, 38, 39, 40, 12, 41, 42, 43, 25, 44, 45, 46].
@@
Another important research direction in modulo ADC is the establishment of theoretical performance guarantees to reconstruct the unfolded signal from modulo samples. In [7], it is shown that the higher-order differences-based reconstruction approach can perfectly reconstruct a finite-energy bandlimited (BL) signal from its modulo samples if the sampling rate used is at least $2\pi e$ ($\approx 17.07$) times the Nyquist rate. This sufficient condition for the sampling rate increases significantly in the presence of bounded noise [7, Theorem 3]. For finite-energy BL signals, perfect reconstruction is possible for any sampling rate approaching the Nyquist rate from above by using a prediction filter with sufficiently long filter length [16]. This approach does not need any side information such as knowledge of unfolded samples and auto-correlation function. However, the proposed prediction filter-based approach did not take into account the impact of finite quantization. Theoretical guarantees for Fourier-based reconstruction had been established in [10]. However, this approach suffers from spectral leakage and may require long observation windows. To avoid spectral leakage, recovery methods based on time-domain thresholding can also be used. Such methods work for generalized modulo sampling models which consider the modulo hysteresis and the folding transients. However, recovery guarantees of thresholding-based recovery fail under ideal modulo nonlinearities. This is explained in Section III-E. The recently developed UNO framework [47] integrates the Unlimited Sampling Framework (USF) [7] with 1-bit quantization, where each folded sample is processed by multiple 1-bit quantizers using varying thresholds. In effect, the set of 1-bit quantizers operates collectively as a multi-bit quantizer. Furthermore, the accompanying theoretical guarantees are expressed through high-probability bounds.
@@
A recent work on modulo ADCs with 1-bit side information [14, 48] established a mean squared error (MSE) performance guarantee when the oversampling factor (OF) is above three and the number of quantization bits is above three. The 1-bit side information identifies the indices in which folding occurred. However, the proofs neglected the spectral leakage effects and the contribution of quantization noise at certain time indices. When the 1-bit folding information is not available in the recovery process, the same MSE guarantees can be achieved by an orthogonal matching pursuit (OMP)-based recovery scheme [48]. However, the sufficient condition for the amplitude quantization bit budget becomes stricter.
In this work, we propose a sliding DFT approach to construct the original input samples from the modulo ADC output samples and an additional 1-bit folding information signal. The recovery procedure uses short-time Fourier transform (STFT) to facilitate successive frame-based unfolding of the modulo samples. To significantly reduce the impact of spectral leakage in STFT computation, a window function with smooth transitions at the edges is multiplied to the frame. A MSE performance guarantee for the proposed recovery method is also established and is compared with the simulated MSE results. Effectively, the use of sliding DFT-based recovery solves the high computational complexity issue of the Fourier-based recovery with 1-bit folding information presented in [14, 48].
@@
  • We develop an unfolding algorithm for modulo ADC output with 1-bit folding information signal based on sliding-window DFT. In contrast to existing Fourier-based reconstruction techniques, the proposed algorithm does not need a long observation window to process the modulo ADC output. This significantly reduces the computational complexity of the recovery algorithm. Furthermore, spectral leakage due to the short frame observations can be reduced by choosing an appropriate window function. The computational complexity of the proposed recovery algorithm is derived in Section III-G.
@@
  • We establish sufficient conditions for the oversampling factor (OF) and quantizer resolution ($b$) of the modulo ADC so that the proposed recovery algorithm can correctly unfold the modulo ADC output. Under these conditions, only the quantization noise contributes to the MSE. The MSE performance guarantee of the sliding DFT-based recovery is derived in Theorem 1.
@@
  • We compare the MSE of the modulo ADC equipped with our proposed recovery procedure to that of the conventional ADC, i.e., no modulo operation, and that of modulo ADC equipped with higher-order differences (HoD)-based recovery method. Under the established sufficient conditions, we show that the MSE of the modulo ADC which uses the sliding DFT recovery algorithm is in $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{3}}\right)$ when spectral leakage is neglected and is in $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{2}}\right)$ otherwise. The parameter $\mathrm{OF}$ is the oversampling factor to be defined in Section II. Meanwhile, the MSE of a conventional ADC is in $\mathcal{O}\left(\frac{1}{\mathrm{OF}}\right)$. We also note that the HoD-based recovery method fails to unfold the modulo ADC output at certain values of $\mathrm{OF}$ when $b=4,5$. Numerical results are presented to demonstrate these behaviors.
@@
Notation: The following notations are used throughout the paper. The sets of real numbers, integers, and natural numbers are denoted as $\mathbb{R}$, $\mathbb{Z}$, and $\mathbb{N}$, respectively. When referring to discrete-time signals, the notation $z[n]$ is used to signify $z(nT_{\mathrm{s}})$, assuming that the sampling period $T_{\mathrm{s}}$ is evident from the context. The first-order difference of a discrete-time signal $x[n]$ is denoted as $\underline{x}[n]=x[n]-x[n-1]$, where x[1]=0x[-1]=0 unless stated otherwise. Vectors and matrices are written in bold format (e.g., $\mathbf{z}$, $\mathbf{A}$) while sets are written in calligraphic format (e.g., $\mathcal{S}$). The cardinality of a set $\mathcal{S}$ is written as $|\mathcal{S}|$. If $\mathbf{x}$ is a vectorized form of some discrete-time signal $x[n]$, then $\underline{\mathbf{x}}$ is the vectorized form of $\underline{x}[n]$. The $\ell_{p}$-norm of a vector $\mathbf{x}$ is written as $\|\mathbf{x}\|_{p}$. The set of values of the discrete-time signal $x[n]$ for $n\in\mathcal{S}$ is denoted as $x_{\mathcal{S}}[n]$. The Moore-Penrose inverse of a matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$ is denoted by $\mathbf{A}^{\dagger}$. We use $\omega$ and $\Omega$ to denote the frequencies for the continuous-time Fourier Transform (CTFT) and discrete-time Fourier Transform (DTFT) spectrum, respectively. We also use the standard Big-O notation $\mathcal{O}(\cdot)$ to describe asymptotic growth rates in this paper.
@@ II-A Signal Acquisition Process
Figure 1: The schematic diagram of the modulo ADC with 1-bit folding information. The 1-bit folding information $c[n]$ is obtained by sampling $c(t)$.
Figure 2: The folding mechanism and folding information generation of the modulo operator block.
We consider the modulo ADC system shown in Figure 1. The input $f(t)$ is a bandlimited signal with (angular) frequency support $\left[-\frac{\omega_{m}}{2},\;+\frac{\omega_{m}}{2}\right]$. This signal is folded by the modulo operator block to produce the modulo signal $f_{\lambda^{\prime}}(t)$. The operation behind the modulo folding mechanism is shown in Figure 2. This feedback structure for the folding operation is adopted from [34]. Comparator 1 triggers a positive value whenever $f_{\lambda^{\prime}}(t)$ crosses $\lambda$ from below. Similarly, comparator 2 triggers a negative value whenever $f_{\lambda^{\prime}}(t)$ crosses $\lambda$ from above. These trigger signals are fed to a discrete voltage generator (DVG). The voltage level of the DVG is increased (resp. decreased) by $2\lambda^{\prime}$ whenever it receives a positive (resp. negative) trigger signal. A detailed circuit-level implementation of the DVG is presented in [34]. Mathematically, the folded signal can be expressed as
@@
To digitize the modulo signal, $f_{\lambda^{\prime}}(t)$ is first sampled at every $T_{s}$ seconds. The angular sampling rate is $\omega_{s}=\frac{2\pi}{T_{s}}=\mathrm{OF}\times\omega_{m}$, where $\mathrm{OF}\geq 1$ is the oversampling factor. We also define $\rho=\frac{1}{\mathrm{OF}}$. Prior to quantization, a (non-subtractive) dither sequence $d[n]$ is added to $f_{\lambda^{\prime}}[n]$. Samples of the dither sequence are drawn i.i.d. from a triangle distribution with amplitude support $\big{(}-\frac{2\lambda}{2^{b}},+\frac{2\lambda}{2^{b}}\big{]}$. The rationale for using triangle dither in the modulo ADC problem setup is explained in Section III-B. The resulting signal is then forwarded to the $b$-bit uniform scalar quantizer $\mathcal{Q}_{b}(\cdot)$ with range $\lambda,+\lambda]$. Consequently, each quantization bin has width $\frac{2\lambda}{2^{b}}$. To prevent the quantizer from being overloaded, the quantizer DR is set to $\lambda=\frac{2^{b}\lambda^{\prime}}{2^{b}-2}$. The quantizer output can be written as
$f_{\lambda^{\prime},\mathrm{q}}[n]=\mathcal{Q}_{b}(f_{\lambda^{\prime}}[n]+d[n])=f[n]+z[n]+\epsilon[n],$ (2)
@@
In addition to $f_{\lambda^{\prime}}(t)$, the modulo operator also generates a signal $c(t)$ according to Figure 2. This signal is a finite rate of innovation (FRI) signal and can be expressed as
$c(t)=$ $g_{2}(t)\ast\left(\sum_{k}g_{1}(t-\tau_{k})\right)=\sum_{k}g(t-\tau_{k}),$
where $\tau_{k}$ is the location of the level crossings, $g_{1}(t)$ is the pulse shape of the trigger signals, $g_{2}(t)$ is the sampling kernel, and $g(t)=(g_{2}\ast g_{1})(t)$. To simplify the analysis, $g(t)$ is assumed to be a rectangular pulse with width $T_{s}$ and unity amplitude. The signal $c(t)$ is sampled at rate $\omega_{s}$ and then fed to a 1-bit comparator to produce $c[n]$. In essence, $c[n]$ provides information if the signal crossed either $\lambda^{\prime}$ or $\lambda^{\prime}$ in the interval $(nT_{\mathrm{s}},(n+1)T_{\mathrm{s}}]$. We note that even though our modulo ADC generates two channel streams, our setup is different from the multichannel modulo sampling structures presented in [23, 24, 25]. The multiple channels in [23, 24, 25] use different modulo thresholds to process the same signal, thereby generating two multi-bit modulo ADC output sample streams. In contrast, our second stream only indicates the position of the folding instances in the first stream. The information carried by this second stream can be represented by a single bit per sample.
Figure 3: Overview of the proposed sliding DFT-based unfolding method for modulo sampling.

### v4

# Short-time Fourier Transform-based Signal Recovery for Modulo Analog-to-Digital Converters

@@ Abstract
This study introduces a short-time Fourier transform–based method for reconstructing signals encoded using modulo analog-to-digital converters with 1-bit folding information. In contrast to existing Fourier-based reconstruction approaches that require complete access to the entire observation, the proposed technique performs reconstruction over short, overlapping segments, enabling significantly lower latency while preserving the recovery accuracy. We also address the spectral leakage introduced by the windowing operation by selecting window parameters that balance the leakage suppression and the computational complexity of the algorithm. In addition, we establish conditions under which the correct unfolding of the modulo samples is guaranteed, leading to a reconstruction error determined solely by the quantization noise at the output. The numerical results demonstrate that the proposed method enables modulo analog-to-digital converters to surpass the mean squared error performance of conventional analog-to-digital converters. Furthermore, the proposed recovery method offers improved reconstruction performance compared with higher-order difference–based recovery, particularly in low-resolution and low–sampling rate regimes.
@@ I Introduction
An analog-to-digital converter (ADC) is a critical element of a data acquisition system because it is responsible for the transformation of continuous-time analog observations from the physical world into a format that can be further processed or stored in digital signal processing (DSP) pipelines. The conversion process in conventional ADCs involves three operations: (1) a sampler that acquires the amplitude of the continuous-time input at uniform time intervals; (2) a quantizer that maps the amplitude samples into a finite set of discrete values; and (3) an encoder that converts the quantized values into binary representations [12]. The sampling rate and quantizer resolution of the ADC were chosen according to a fidelity criterion and a power consumption budget. Theoretically, the sampling rate requirement for alias-free reconstruction and the quantization resolution needed to achieve some distortion level are governed by the Shannon-Nyquist sampling theorem [48] and the rate-distortion theory [3], respectively. However, increasing these parameters also incurs a substantial increase in the ADC power consumption. Specifically, the ADC power consumption scales exponentially with the number of quantization bits and linearly with the sampling rate [51].
@@
Another important parameter that affects the performance of an ADC is its dynamic range (DR). Owing to the finite number of quantization bits in an ADC, a trade-off exists between its dynamic range and resolution. An input signal may drive the ADC to saturation if the DR of the input signal exceeds the DR of the ADC. This issue can be mitigated by increasing the DR of the ADC. However, doing so under a fixed quantization bit budget increases the quantization noise power, causing weak input signals to be overwhelmed by the quantization noise. The presence of both weak and strong components in the input signal increases the ADC resolution required to achieve a specific fidelity criterion [31].
@@
A promising solution to address the DR bottleneck of ADCs is the modulo sampling framework [7, 9]. The idea behind modulo sampling is to apply a nonlinear folding operation called modulo to the input signal before the ADC. This modulo pre-processing stage ensures that the signal range is confined within the DR of the ADC. Moreover, for a given amplitude quantization bit budget, modulo ADCs can achieve better digital resolution than conventional ADCs as demonstrated in previous works [57, 27].
@@
Because the modulo operation intentionally introduces nonlinear distortion to the signal, several studies on modulo sampling have focused on the development of robust recovery techniques to correct the distortion induced by the modulo operation. Reconstruction techniques for modulo sampling include higher-order differences-based approaches [9], discrete Fourier transform (DFT)-based techniques [6, 1, 53, 45, 4, 2], prediction-based filtering [43, 38], iterative methods [26], sparse signal modulo recovery [8, 37, 47, 41], multi-channel modulo structures [21, 23, 49, 20], and thresholding-based schemes[18, 19, 16, 17, 29]. There have also been research efforts geared toward the practical hardware implementation of modulo ADCs [6, 30, 13, 36, 56, 46] and their potential applications in radar systems, communication systems, imaging, and electroencephalogram (EEG) recovery [10, 39, 14, 15, 53, 50, 22, 32, 49, 33, 54, 35, 52].
@@
Another important direction in modulo ADC research is the establishment of theoretical performance guarantees to reconstruct the unfolded signal from the modulo samples. In [9], it was shown that the higher-order differences-based reconstruction approach can perfectly reconstruct a finite-energy bandlimited (BL) signal from its modulo samples if the sampling rate used is at least $2\pi e$ ($\approx 17.07$) times the Nyquist rate. This sufficient condition for the sampling rate significantly increases in the presence of bounded noise [9, Theorem 3]. For finite-energy BL signals, perfect reconstruction is possible for any sampling rate approaching the Nyquist rate from above by using a prediction filter with a sufficiently long filter length [43]. This approach does not require any side information such as knowledge of unfolded samples and the auto-correlation function. However, the proposed prediction filter-based approach does not take into account the impact of finite quantization. Theoretical guarantees for the Fourier-based reconstruction were established in [6]. However, this approach suffers from spectral leakage and may require long observation windows. To avoid spectral leakage, recovery methods based on time-domain thresholding can be used [18, 19, 16, 17, 29]. Such methods work for generalized modulo sampling models that consider modulo hysteresis and folding transients. However, the recovery guarantees established for thresholding-based recovery methods fail under ideal modulo nonlinearities, as explained in Section III-E. The recently developed UNO framework [11] integrates the Unlimited Sampling Framework (USF) [9] with 1-bit quantization, where each folded sample is processed by multiple 1-bit quantizers using varying thresholds. In effect, the set of 1-bit quantizers operates collectively as a multi-bit quantizer. Furthermore, the accompanying theoretical guarantees are expressed through high-probability bounds.
@@
A recent line of research on modulo ADCs with 1-bit folding side information [4, 5] established a mean squared error (MSE) performance guarantee when the oversampling factor (OF) exceeds three and the quantizer resolution exceeds three bits. The 1-bit side information identifies the time indices at which folding occurs. However, the analysis in these studies neglects both spectral leakage effects and the contribution of quantization noise at certain time indices. When the 1-bit folding side information is unavailable, the same MSE guarantees can be recovered using an orthogonal matching pursuit (OMP)–based reconstruction scheme [5], although this requires a stricter condition on the amplitude quantization bit budget. Additionally, the Fourier-based recovery methods in [4, 5] incur very high computational complexity, rendering them impractical for very long sequences.
In this study, we propose a short-time Fourier transform (STFT)–based recovery method for reconstructing the original input samples from the output of a modulo ADC equipped with 1-bit folding side information. The method performs unfolding on short, overlapping frames of the signal, enabling successive, low-latency reconstruction rather than requiring access to the entire sequence at once. We employ a smooth window function with controlled roll-off factor to mitigate the spectral leakage introduced by the STFT windowing operation. We further established a mean-squared error (MSE) performance guarantee for the proposed method and validated the analysis through numerical experiments. This STFT-based recovery framework substantially reduces the computational complexity relative to the Fourier-based recovery schemes, especially for long sequences. Moreover, we provide a more comprehensive analysis of the MSE guarantees by considering the contribution of the spectral leakage effects inherent in Fourier-based recovery methods. This has not been accounted for in the earlier works.
Figure 1: Schematic diagram of the modulo ADC with 1-bit folding information. The input signal is wrapped by the modulo operation whenever it exceeds the converter’s dynamic range, producing a folded output sequence. A folding detector simultaneously generates a continuous-time folding signal $c(t)$, and the corresponding discrete-time folding information $c[n]$ is obtained by sampling this signal.
Figure 2: Modulo operator mechanism to generate the folded signal $f_{\lambda^{\prime}}(t)$ and the folding information signal $c(t)$. The comparators detect folding events and provide control signals to the discrete voltage generator that produces the wrapped output. At the same time, comparator outputs are processed to generate the continuous-time folding information $c(t)$.
@@
  • We develop an STFT-based recovery algorithm to reconstruct the input signal from the output of a modulo ADC equipped with 1-bit folding information. Unlike existing Fourier-based reconstruction techniques, the proposed method does not require the processing of a long observation window or the entire data sequence before producing unfolded samples, thereby substantially reducing the computational complexity. Moreover, the spectral leakage introduced by short-frame processing is mitigated through the use of a smoother window function with gradual transitions at the boundaries. The STFT window design considerations and computational complexity of the proposed algorithm are analyzed in Sections II-D and III-G, respectively.
@@
  • We derive sufficient conditions for the oversampling factor ($\mathrm{OF}$) and quantizer resolution $\mathrm{b}$ under which the proposed STFT-based algorithm reliably unfolds the modulo ADC output. When these conditions are met, the only source of the reconstruction error is the quantization noise. The MSE performance guarantee is formalized in Theorem 1.
@@
  • We compare the MSE of a modulo ADC equipped with the proposed STFT-based recovery algorithm against that of (i) a conventional ADC without modulo operation and (ii) a modulo ADC using the higher-order differences (HoD)-based recovery method. Under the established sufficient conditions, we show that the MSE of the STFT-based recovery scales as $\mathcal{O}\!\left(1/\mathrm{OF}^{3}\right)$ when spectral leakage is negligible, and as $\mathcal{O}\!\left(1/\mathrm{OF}^{2}\right)$ when leakage is present. Here, $\mathrm{OF}$ denotes the oversampling factor as defined in Section II. In contrast, the MSE of a conventional ADC scales only as $\mathcal{O}\!\left(1/\mathrm{OF}\right)$. We further observed that the HoD-based recovery method fails to reliably unfold the modulo ADC output at certain oversampling factors when $b=4$ or $5$. Numerical results are provided to illustrate these behaviors and validate the theoretical predictions.
The remainder of the paper is organized as follows: Section II describes the system model and the proposed reconstruction algorithm. Section III presents the theoretical guarantees of the proposed recovery algorithm in terms of recovery performance and computational complexity. Section IV provides the numerical results that validates the theoretical analysis of the proposed method. Section V concludes the paper.
@@
Notation: The following notations are used throughout this paper. The sets of real numbers, integers, and natural numbers are denoted by $\mathbb{R}$, $\mathbb{Z}$, and $\mathbb{N}$, respectively. When referring to discrete-time signals, the notation $z[n]$ is used to signify $z(nT_{\mathrm{s}})$, assuming that the sampling period $T_{\mathrm{s}}$ is evident in the context. The first-order difference of a discrete-time signal $x[n]$ is denoted by $\underline{x}[n]=x[n]-x[n-1]$, where x[1]=0x[-1]=0 unless stated otherwise. Vectors and matrices are written in bold format (e.g., $\mathbf{z}$, $\mathbf{A}$) while sets are written in calligraphic format (e.g., $\mathcal{S}$). The cardinality of a set $\mathcal{S}$ is written as $|\mathcal{S}|$. If $\mathbf{x}$ is a vectorized form of some discrete-time signal $x[n]$, then $\underline{\mathbf{x}}$ is the vectorized form of $\underline{x}[n]$. The $\ell_{p}$-norm of a vector $\mathbf{x}$ is written as $\|\mathbf{x}\|_{p}$. The set of values of the discrete-time signal $x[n]$ for $n\in\mathcal{S}$ is denoted as $x_{\mathcal{S}}[n]$. The Moore-Penrose inverse of the matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$ is denoted by $\mathbf{A}^{\dagger}$. We use $\omega$ and $\Omega$ to denote the frequencies for the continuous-time Fourier Transform (CTFT) and discrete-time Fourier Transform (DTFT) spectra, respectively. We also use the standard Big-O notation $\mathcal{O}(\cdot)$ to describe the asymptotic growth rates in this paper.
@@ II System Model and Reconstruction Algorithm
In this section, we describe the signal acquisition process of the modulo ADC and introduce the proposed short-time Fourier Transform (STFT)-based recovery framework. We first present the hardware model and mathematical formulation of the modulo folding operation, including the generation of the 1-bit folding information that indicates the folding instances. We also outline the proposed STFT-based signal recovery scheme and highlight its advantages over existing Fourier-based recovery methods. We also provide design guidelines for selecting the STFT window parameters.
@@ II-A Signal Acquisition Process
We consider the modulo ADC system shown in Figure 1. The input $f(t)$ is a bandlimited signal with (angular) frequency support $\left[-\frac{\omega_{m}}{2},\;+\frac{\omega_{m}}{2}\right]$. This signal is folded by the modulo operator block to produce the modulo signal $f_{\lambda^{\prime}}(t)$. The operation behind the modulo folding mechanism is illustrated in Figure 2. This feedback structure for the folding operation is adopted from [36]. Comparator 1 triggers a positive output whenever $f_{\lambda^{\prime}}(t)$ crosses $\lambda$ from below. Similarly, comparator 2 triggers a negative output whenever $f_{\lambda^{\prime}}(t)$ crosses $\lambda$ from above. These trigger signals are fed to a discrete voltage generator (DVG). The voltage level of the DVG is increased (resp. decreased) by $2\lambda^{\prime}$ whenever it receives a positive (resp. negative) trigger signal. A detailed circuit-level implementation of the DVG is presented in [36]. Mathematically, the folded signal can be expressed as
@@
To digitize the modulo signal, $f_{\lambda^{\prime}}(t)$ is first sampled at every $T_{s}$ seconds. The angular sampling rate is $\omega_{s}=\frac{2\pi}{T_{s}}=\mathrm{OF}\times\omega_{m}$, where $\mathrm{OF}\geq 1$ is the oversampling factor. We also define $\rho=\frac{1}{\mathrm{OF}}$. Prior to quantization, a (non-subtractive) dither sequence $d[n]$ is added to $f_{\lambda^{\prime}}[n]$. Samples of the dither sequence are drawn i.i.d. from a triangle distribution with amplitude support $\big(-\frac{2\lambda}{2^{b}},+\frac{2\lambda}{2^{b}}\big]$. The rationale for using triangle dither in the modulo ADC problem setup is explained in Section III-B. The resulting signal is then forwarded to the $b$-bit uniform scalar quantizer $\mathcal{Q}_{b}(\cdot)$ with range $\lambda,+\lambda]$. Consequently, each quantization bin has width $\frac{2\lambda}{2^{b}}$. To prevent the quantizer from being overloaded, the quantizer DR is set to $\lambda=\frac{2^{b}\lambda^{\prime}}{2^{b}-2}$.
The quantizer output can be written as
$f_{\lambda^{\prime},\mathrm{q}}[n]=$ $\mathcal{Q}_{b}(f_{\lambda^{\prime}}[n]+d[n])$
$=$ $f[n]+z[n]+\epsilon[n],$ (2)
@@
Figure 3: Overview of the proposed STFT-based recovery method for modulo sampling. The proposed recovery method is composed of three steps: (1) out-of-band DFT computation, (2) estimation of the modulo residue via 1-bit folding information and out-of-band DFT values, and (3) removal of modulo residue from folded signal.
In addition to $f_{\lambda^{\prime}}(t)$, the modulo operator also generates a signal $c(t)$ as shown in Figure 2. This signal is a finite rate of innovation (FRI) signal and can be expressed as
$c(t)=$ $g_{2}(t)\ast\left(\sum_{k}g_{1}(t-\tau_{k})\right)$
$=$ $\sum_{k}g(t-\tau_{k}),$ (3)
where $\tau_{k}$ is the location of the level crossings, $g_{1}(t)$ is the pulse shape of the trigger signals, $g_{2}(t)$ is the sampling kernel, and $g(t)=(g_{2}\ast g_{1})(t)$. To simplify the analysis, $g(t)$ is assumed to be a rectangular pulse with width $T_{s}$ and unity amplitude. The signal $c(t)$ is sampled at rate $\omega_{s}$ and then fed to a 1-bit comparator to produce $c[n]$. In essence, $c[n]$ provides information on whether the signal crosses either $\lambda^{\prime}$ or $\lambda^{\prime}$ in the interval $(nT_{\mathrm{s}},(n+1)T_{\mathrm{s}}]$. We note that even though our modulo ADC generates two channel streams, our setup is different from the multi-channel modulo sampling structures presented in [21, 23, 49]. The multiple channels in [21, 23, 49] use different modulo thresholds to process the same signal, thereby generating two multi-bit modulo ADC output sample streams. In contrast, our second stream only indicates the position of the folding instances in the first stream. The information carried by this second stream can be represented as a single bit per sample.